### v1

# Some Remarks on Positive/Negative Feedback

Abstract
In the context of unstable systems with control, a commonly-held precept is that negative and positive feedback cannot both be stabilizing. The canonical linear prototype is the scalar system $\dot{x}=u$ which, under negative linear feedback u=kxu=-kx ($k>0$) is exponentially stable for all $k>0$, whereas the inherent lack of exponential instability of the uncontrolled system is amplified by positive feedback $u=kx$ ($k>0)$. By contrast, for nonlinear systems it is shown that this intuitively-appealing dichotomy may fail to hold.
keywords:
Nonlinear control systems; positive and negative feedback.
fundingfundingfootnotetext: T. Berger acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 544702565.
1 Nonlinear scalar system
Consider a scalar system, with state $x$, control $u$, and (unknown) bounded perturbation $p$, of the form
$\dot{x}(t)=f(p(t),x(t))+g(u(t)),\quad x(0)=x^{0}\in\mathbb{R},$ (1.1)
where $f\colon\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, $g\colon\mathbb{R}\to\mathbb{R}$, and $p\colon\mathbb{R}_{\geq 0}\to\mathbb{R}$ are continuous. The control objective is to determine a continuous feedback function $h\colon\mathbb{R}_{\geq 0}\times\mathbb{R}\to\mathbb{R}$, with an associated parameter $\eta\in\{-1,+1\}$, such that, for every $x^{0}\in\mathbb{R}$ and bounded $p\in\mathcal{C}(\mathbb{R}_{\geq 0},\mathbb{R})$, application of the feedback $u(t)=-\eta h(t,x(t))$ to (1.1) yields a feedback-controlled initial-value problem
$\dot{x}(t)=f(p(t),x(t))+g(-\eta h(t,x(t)),\quad x(0)=x^{0}$
has a solution, every solution has a global extension to a solution on $\mathbb{R}_{\geq 0}$, every global solution is such that $x(t)\to 0$ as $t\to\infty$, and the control input $u\colon t\mapsto-\eta h(t,x(t))$ is bounded.
If $\eta=+1$ (respectively, $\eta=-1$), then the feedback $u(t)=-\eta h(t,x(t))$ is deemed to be negative (respectively, positive). If negative feedback ensures benign behaviour in the form of convergence of the state to zero, then intuition might suggest that positive feedback causes malign behaviour in the form of non-convergence of the state to zero.
This intuition finds voice in, for example, [1]: Feedback can upregulate or downregulate a process. We refer to upregulation as positive feedback and downregulation as negative feedback. Positive feedback … acts as a destabilizing mechanism for the process. Negative feedback … acts as a stabilizing mechanism for the process.
We show, by a simple example, that this positive/negative feedback dichotomy can be invalid. In particular, we construct an example (with arbitrary $f$ and $p$) which, under either negative or positive feedback, ensures convergence of the state to zero whilst maintaining boundedness of the control input.
2 Feedback function
Let $\varphi\colon\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ be a continuously differentiable bijection with the property that, for some $c_{\varphi}>0$, $\dot{\varphi}(t)\leq c_{\varphi}\big(1+\varphi(t)\big)$ for all $t\geq 0$ (the simplest example being the identity function with $c_{\varphi}=1$). Define the set
$\mathcal{F}:=\{(t,\xi)\in\mathbb{R}_{\geq 0}\times\mathbb{R}\mid~\varphi(t)|\xi|<1\},$ (2.1)
the continuous function
$h\colon\mathcal{F}\to\mathbb{R},~~(t,\xi)\mapsto\alpha(\varphi(t)|\xi|)\varphi(t)\xi,$ (2.2)
$\text{with}~\alpha\colon[0,1)\to[1,\infty),~s\mapsto\frac{1}{1-s},$
and the feedback-controlled initial-value problem, on the domain $\mathcal{F}$,
$\dot{x}(t)=f(p(t),x(t))+g(-\eta h(t,x(t)),~~x(0)=x^{0}$ (2.3)
with $p\in\mathcal{C}(\mathbb{R}_{\geq 0},\mathbb{R})$. By a solution we mean a continuous function $x$ on a left-closed interval II (with left endpoint 0) and $\text{graph}(x)\subset\mathcal{F}$, satisfying (2.3). A solution is maximal if it has no proper right extension that is also a solution. By the standard theory of ordinary differential equations, for each $x^{0}\in\mathbb{R}$, there exists a solution, and every solution has a maximal extension; moreover, the closure of graph of a maximal solution is not a compact subset of $\mathcal{F}$.
[…]
Figure 1: Domain $\mathcal{F}$.
Loosely speaking, the role of the feedback is to maintain the system’s evolution away from the boundary of the domain $\mathcal{F}$ (wherein singularity resides), in which case the evolution continues indefinitely to the right, with transient and asymptotic behaviour determined by the choice of $\varphi$.
3 Efficacy of both negative and positive feedback
Let $g$ be given by
$g\colon\mathbb{R}\to\mathbb{R},~~v\mapsto v\sin\big(\ln(1+|v|)\big).$ (3.1)
The following technicality will play a central role.
Proposition 3.1.
Let $f\colon\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ be continuous, $g$ as in (3.1), and $K,P\subset\mathbb{R}$ compact. Define $V:=[-1,-\tfrac{1}{2}]\cup[\tfrac{1}{2},1]$. The function
$\chi\colon$ $\mathbb{R}\to\mathbb{R},$
$s\mapsto\min\{vf(\rho,\xi)+vg(sv)\mid\rho\in P,~\xi\in K,~v\in V\}$
is continuous and is such that, for each $\eta\in\{-1,+1\}$, $\sup_{s\geq 0}\chi(\eta s)=\infty$.
Proof.
For notational convenience, introduce the function
$F\colon\mathbb{R}^{3}\times\mathbb{R}\to\mathbb{R},~~(z,s)=\big((\rho,\xi,v),s\big)\mapsto vf(\rho,\xi)+vg(sv)$
and write $Z:=P\times K\times V$, a compact set. Observe that
$\chi\colon s\mapsto\min_{z\in Z}F(z,s)$
and so, for each $s\in\mathbb{R}$, there exists $z_{s}\in Z$ such that $\chi(s)=F(z_{s},s)$. Fix $s\in\mathbb{R}$ arbitrarily. We will show that $\chi$ is continuous at $s$. Let $\varepsilon>0$. By continuity of $F$, for each $z\in Z$, there exists $\delta(z)>0$ such that, for all $\zeta,\sigma\in\mathbb{R}$ we have
$\|\zeta-z\|+|\sigma-s|<\delta(z)\implies|F(\zeta,\sigma)-F(z,s)|<\varepsilon$
and so $\cup_{z\in Z}\{(\zeta,\sigma)\mid\|\zeta-z\|+|\sigma-s|<\delta(z)\}$ is an open cover of $Z\times\{s\}$. By compactness of the latter, there exists a finite sub-cover. In particular, there exist finitely many points $z_{1},\ldots,z_{n}\in Z$ such that
$Z\times\{s\}\subset\cup_{i=1}^{n}\{(\zeta,\sigma)\mid\|\zeta-z_{i}\|+|\sigma-s|<\delta(z_{i})\}.$
Define $\delta^{*}:=\min\{\delta(z_{1}),\ldots,\delta(z_{n})\}$ and observe that
$\forall\,\zeta,\sigma\in\mathbb{R}\colon~~z\in Z,~|\sigma-s|<\delta^{*}\implies|F(z,\sigma)-F(z,s)|<\varepsilon.$
It follows that, for all $\sigma\in\mathbb{R}$ with $|\sigma-s|<\delta^{*}$ we have
$\varepsilon>|F(z_{s},\sigma)-F(z_{s},s)|\geq F(z_{s},\sigma)-F(z_{s},s)\geq\chi(\sigma)-\chi(s)$
and, similarly,
$\varepsilon>|F(z_{\sigma},\sigma)-F(z_{\sigma},s)|\geq F(z_{\sigma},s)-F(z_{\sigma},\sigma)\geq\chi(s)-\chi(\sigma).$
Therefore, $|\chi(\sigma)-\chi(s)|<\varepsilon$ for all $\sigma\in\mathbb{R}$ with $|\sigma-s|<\delta^{*}$, and so $\chi$ is continuous at $s$.
It remains to show that, for each $\eta\in\{-1,+1\}$, $\sup_{s\geq 0}\chi(\eta s)=+\infty$. Define the increasing and unbounded sequence $(s_{n})$ by
$s_{n}:=\tfrac{1}{2}e^{(n+1)\pi}-1>0\quad\forall\,n\in\mathbb{N},$
whence
$\ln(1+s_{n})=(n+1)\pi-\ln 2\quad\forall\,n\in\mathbb{N}.$
Also, invoking the inequality $e^{\pi/2}>4$,
$1+\tfrac{1}{2}s_{n}=\tfrac{1}{2}+\tfrac{1}{4}e^{(n+1)\pi}>e^{n\pi}(\tfrac{1}{4}e^{\pi})>e^{(n+\frac{1}{2})\pi}$
and so
$\ln(1+\tfrac{1}{2}s_{n})>n\pi+\tfrac{\pi}{2}\quad\forall\,n\in\mathbb{N}.$
We may now infer
$n\pi+\tfrac{\pi}{2}<\ln(1+s_{n}|v|)\leq(n+1)\pi-\ln 2\quad\forall\,v\in V~~\forall\,n\in\mathbb{N}.$
First consider the case of $n$ even and write $n=2k$. Since the function $\sin$ is decreasing on the interval $\big[\tfrac{\pi}{2},\pi-\ln 2\big]$, for all $v\in V$ and all $k\in\mathbb{N}$ we have
$1$ $=\sin(2k\pi+\tfrac{\pi}{2})>\sin(\ln(1+s_{2k}|v|))$
$\geq\sin((2k+1)\pi-\ln 2)=\sin(\pi-\ln 2)=\sin(\ln 2)>0,$
whence
$vg(s_{2k}v)=s_{2k}v^{2}\sin(\ln(1+s_{2k}|v|))\geq\tfrac{1}{4}s_{2k}\sin(\ln 2).$
Therefore, writing
$c_{1}:=\min\{vf(\rho,\xi)\mid\rho\in P,~\xi\in K,~v\in V\},$
we have
$\chi(s_{2k})\geq c_{1}+\min\{vg(s_{2k}v)\mid v\in V\}\geq c_{1}+\tfrac{1}{4}s_{2k}\sin(\ln 2)$
for all $k\in\mathbb{N}$ and so $\sup_{s\geq 0}\chi(s)=+\infty$.
Now consider the case of $n$ odd and write n=2k1n=2k-1. Since the function $\sin$ is increasing on the interval $\big[-\tfrac{\pi}{2},-\ln 2\big]$, we have
$-1$ $=\sin\big((2k-1)\pi+\tfrac{\pi}{2}\big)<\sin(\ln(1+s_{2k-1}|v|))$
$\leq\sin(2k\pi-\ln 2)=-\sin(\ln 2)\quad\forall\,v\in V,~~\forall\,k\in\mathbb{N},$
whence
$vg(-s_{2k-1}v)$ $=-s_{2k-1}v^{2}\sin(\ln(1+s_{2k-1}|v|))$
$\geq\tfrac{1}{4}s_{2k-1}\sin(\ln 2)\qquad\forall\,v\in V,~~\forall\,k\in\mathbb{N}.$
Therefore,
$\chi(-s_{2k-1})>c_{1}+\tfrac{1}{4}s_{2k-1}\sin(\ln 2)\quad\forall\,k\in\mathbb{N}$
and so $\sup_{s\geq 0}\chi(-s)=+\infty$. ∎
We arrive at the main focus of the note.
Proposition 3.2.
Let $f\in\mathcal{C}(\mathbb{R},\mathbb{R})$ be arbitrary, $g$ as in (3.1), and $h$ as in (2.2). For each $\eta\in\{-1,+1\}$, every $x^{0}\in\mathbb{R}$, and every bounded $p\in\mathcal{C}(\mathbb{R}_{\geq 0},\mathbb{R})$, (2.3) has a solution and every solution has a global extension. Every global solution $x\colon\mathbb{R}_{\geq 0}\to\mathbb{R}$ is such that $x(t)\to 0$ as $t\to\infty$, and the control function $u\colon t\mapsto-\eta h(t,x(t))$ is bounded.
Proof.
Let $\eta\in\{-1,+1\}$, $x^{0}\in\mathbb{R}$ and bounded $p\in\mathcal{C}(\mathbb{R}_{\geq 0},\mathbb{R})$ be arbitrary. Let compact $P\subset\mathbb{R}$ be such that $p(t)\in P$ for all $t\geq 0$. As already noted, by the standard theory of ordinary differential equations, we know that (2.3) has a solution and every solution can be maximally extended. Let $x\colon[0,\omega)\to\mathbb{R}$ be a maximal solution. It suffices to show existence of  $\varepsilon>0$ such that $\varphi(t)|x(t)|<1-\varepsilon$ for all $t\in[0,\omega)$, in which case it immediately follows that $\omega=\infty$ (as, otherwise, $\text{graph}(x)$ has compact closure in $\mathcal{F}$ which is impossible), $x(t)\to 0$ as $t\to\infty$ (since $\varphi$ is unbounded), and $|u(t)|=|h(t,x(t))|\leq(1-\varepsilon)\alpha(1-\varepsilon)<\infty$ for all $t\geq 0$. The task of establishing the existence of such  $\varepsilon>0$ is, in turn, equivalent to proving boundedness of the function $k\colon t\mapsto\alpha(\varphi(t)|x(t)|)$. This we proceed to show.
Choose $\tau\in(0,\omega)$ arbitrarily, fix $\mu>1/\varphi(\tau)$ such that $\mu\geq\max_{t\in[0,\tau]}|x(t)|$, and write $\varphi_{\mu}\colon t\mapsto\max\{\mu^{-1},\varphi(t)\}$. Since $\text{graph}(x)\in\mathcal{F}$ and $\varphi$ (bijective) is strictly increasing, we may infer that
$|x(t)|\leq\frac{1}{\varphi_{\mu}(t)}\quad\forall\,t\in[0,\omega)$
and so, a fortiori, $x(t)\in K:=[-\mu,\mu]$ for all $t\in[0,\omega)$. As before, define
$\chi\colon$ $\mathbb{R}_{\geq 0}\to\mathbb{R},$
$s\mapsto\min\{vf(\rho,\xi)+vg(sv)\mid\rho\in P,~\xi\in K,~v\in V\},$
$\qquad\qquad\qquad\qquad\qquad\qquad V:=[-1,-\tfrac{1}{2}]\cup[\tfrac{1}{2},1].$
Let $\nu$ be the function
$\nu\colon~~\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0},~~s\mapsto\nu(s):=\max\{\chi(\eta s),0\}.$
By Proposition 3.1, $\chi$ is continuous with $\sup_{s\geq 0}\chi(\eta s)=+\infty$ and so $\nu$ is continuous and surjective. Fix $\kappa_{0}>\alpha(1/2)$ arbitrarily and define an increasing sequence $(\kappa_{n})$ in $(\kappa_{0},\infty)$ as follows: choose $\kappa_{1}>\kappa_{0}$ such that $\nu(\kappa_{1})=\kappa_{0}+1$ and set
$\kappa_{n}:=\inf\{\kappa>\kappa_{n-1}\mid\nu(\kappa)=\kappa_{0}+n\}\quad\forall\,n\geq 2.$
$\kappa$$\nu(\kappa)$$\kappa_{0}+1$$\kappa_{0}+2$$\kappa_{0}+3$$\kappa_{1}$$\kappa_{2}$$\kappa_{3}$
Figure 2: Schematic construction of the sequence $(\kappa_{n})$.
Now, we address the question of boundedness of $k(\cdot)$.
Seeking a contradiction, suppose that $k(\cdot)$ is unbounded. This supposition, together with the observation that $k(0)=1<\kappa_{0}$, ensures that the following is a well-defined sequence  $(\tau_{n})$ in $(0,\omega)$
$\tau_{n}:=\inf\{t\in(0,\omega)\mid k(t)=\kappa_{n}\},\quad n\in\mathbb{N}.$
with the properties
$\forall\,n\in\mathbb{N}:\ \tau_{n}<\tau_{n+1},\quad k(\tau_{n})=\kappa_{n},\quad\nu(k(\tau_{n}))=\kappa_{0}+n.$
[…]
Figure 3: Schematic construction of the sequence $(\sigma_{n})$.
Now define the sequence $(\sigma_{n})$ in $(0,\omega)$ by
$\sigma_{n}:=\sup\{t\in[\tau_{n},\tau_{n+1}]\,\mid\nu(k(t))=\kappa_{0}+n\},\quad n\in\mathbb{N}.$
Observe that $\sigma_{n}<\tau_{n+1}$, $k(\sigma_{n})<k(\tau_{n+1})$ and
$\kappa_{0}+n<\nu(k(t))=\chi(\eta k(t))\leq\nu(k(\tau_{n+1}))=\kappa_{0}+n+1$
for all $t\in(\sigma_{n},\tau_{n+1}]$ and all $n\in\mathbb{N}$. Let $n\in\mathbb{N}$ be arbitrary and suppose that, for some $t\in[\sigma_{n},\tau_{n+1}]$, $\varphi(t)x(t)\not\in V$. Then $\varphi(t)|x(t)|<1/2$ and
$\alpha(\varphi(t)|x(t)|)=k(t)<\alpha(1/2)<\kappa_{0}<\kappa_{n}<\kappa_{n+1}=k(\tau_{n+1})$
Therefore, there exists $t^{*}\in(\sigma_{n},\tau_{n+1})$ such that $k(t^{*})=\kappa_{n}$ whence the contradiction
$\kappa_{0}+n<\nu(k(t^{*}))=\nu(\kappa_{n})=\kappa_{0}+n.$
Thus, we have shown that
$\varphi(t)x(t)\in V\quad\forall\,t\in[\sigma_{n},\tau_{n+1}]\quad\forall\,n\in\mathbb{N}.$
By symmetry of $V$, we know that $\varphi(t)x(t)\in V$ if, and only if, $\varphi(t)x(t)\in V$, and so we may infer that
$-\varphi(t)x(t)f(p(t),x(t))-\varphi(t)x(t)g(-\eta k(t)\varphi(t)x(t))$
$\ \geq\min\{vf(\rho,\xi)+vg(\eta k(t)v)\mid\rho\in P,~\xi\in K,~v\in V\}$
$\ =\chi(\eta k(t))\quad\forall\,t\in[\sigma_{n},\tau_{n+1}]\quad\forall\,n\in\mathbb{N}$
Also (noting that, by bijectivity, $\varphi$ is strictly increasing), we have
$\varphi(t)\dot{\varphi}(t)x(t)^{2}$ $\leq\frac{\dot{\varphi}(t)}{\varphi(t)}\leq c_{\varphi}\big((1/\varphi(t))+1\big)$
$\leq c_{\varphi}\big((1/\varphi(\tau_{1}))+1\big)=:\gamma\qquad\forall\,t\in[\tau_{1},\omega).$
We may now deduce that
$\tfrac{1}{2}\tfrac{\mathrm{d}}{\mathrm{d}t}(\varphi(t)x(t))^{2}$ $=\varphi(t)\dot{\varphi}(t)x(t)^{2}$
$\quad+\varphi(t)^{2}x(t)\big(f(p(t),x(t))+g(-\eta h(t,x(t))\big)$
$\leq\gamma-\varphi(t)\big(-\varphi(t)x(t)f(p(t),x(t))$
$\quad-\varphi(t)x(t)g(-\eta k(t)\varphi(t)x(t))\big)$
$\leq\gamma-\varphi(t)\chi(\eta k(t))$
$\leq\gamma-\varphi(\tau_{1})\nu(k(t))\leq\gamma-(\kappa_{0}+n)\varphi(\tau_{1})$
for all $t\in[\sigma_{n},\tau_{n+1}]$ and all $n\in\mathbb{N}$. Choose $n\in\mathbb{N}$ sufficiently large so that $\gamma-(\kappa_{0}+n)\varphi(\tau_{1})<0$, in which case
$\big(\varphi(\tau_{n+1})x(\tau_{n+1})\big)^{2}<\big(\varphi(\sigma_{n})x(\sigma_{n})\big)^{2}$
and we arrive at a contradiction
$k(\sigma_{n})$ $<k(\tau_{n+1})=\alpha(\varphi(\tau_{n+1})|x(\tau_{n+1})|)$
$<\alpha(\varphi(\sigma_{n})|x(\sigma_{n})|)=k(\sigma_{n}).$
Therefore, the supposition of unboundedness of $k(\cdot)$ is false. ∎
4 Conclusion
The claim in this note is purely existential in nature in the sense that a commonly-perceived dichotomy of negative and positive feedback is shown not to be valid in general. This has been done by proving the existence of a nonlinear counter-example: the practical utility and quantitative behaviour of the feedback employed therein has not been addressed. Whilst the counter-example is open to a criticism of artificiality, it does, however, serve to illustrate that the dichotomy “negative/positive feedback” equates to “good/bad behaviour” may be inaccurate. The counter-example is based on a wider body of work in the area of “funnel control” (a terminology consistent with Figure 1). The interested reader is referred to [2] for a review of this area.

### v2

@@
 Abstract
-In the context of unstable systems with control, a commonly-held precept is that negative and positive feedback cannot both be stabilizing.
-The canonical linear prototype is the scalar system $\dot{x}=u$ which, under negative linear feedback u=kxu=-kx ($k>0$) is exponentially stable for all $k>0$, whereas the inherent lack of exponential instability of the uncontrolled system is amplified by positive feedback $u=kx$ ($k>0)$.
-By contrast, for nonlinear systems it is shown that this intuitively-appealing dichotomy may fail to hold.
+In the context of linear control systems, a commonly-held intuition is that negative and positive feedback cannot both be stability enhancing.
+The canonical linear prototype is the scalar system $\dot{x}=u$ which, under negative linear feedback u=kxu=-kx ($k>0$) is exponentially stable for all $k>0$, whereas the lack of exponential instability of the (marginally stable) uncontrolled system is amplified by positive feedback $u=kx$ ($k>0)$.
+By contrast, for nonlinear systems it is shown, by example, that this intuitive dichotomy may fail to hold.
 keywords:
@@
 Consider a scalar system, with state $x$, control $u$, and (unknown) bounded perturbation $p$, of the form
-$\dot{x}(t)=f(p(t),x(t))+g(u(t)),\quad x(0)=x^{0}\in\mathbb{R},$ (1.1)
-where $f\colon\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, $g\colon\mathbb{R}\to\mathbb{R}$, and $p\colon\mathbb{R}_{\geq 0}\to\mathbb{R}$ are continuous.
-The control objective is to determine a continuous feedback function $h\colon\mathbb{R}_{\geq 0}\times\mathbb{R}\to\mathbb{R}$, with an associated parameter $\eta\in\{-1,+1\}$, such that, for every $x^{0}\in\mathbb{R}$ and bounded $p\in\mathcal{C}(\mathbb{R}_{\geq 0},\mathbb{R})$, application of the feedback $u(t)=-\eta h(t,x(t))$ to (1.1) yields a feedback-controlled initial-value problem
-$\dot{x}(t)=f(p(t),x(t))+g(-\eta h(t,x(t)),\quad x(0)=x^{0}$
-has a solution, every solution has a global extension to a solution on $\mathbb{R}_{\geq 0}$, every global solution is such that $x(t)\to 0$ as $t\to\infty$, and the control input $u\colon t\mapsto-\eta h(t,x(t))$ is bounded.
-If $\eta=+1$ (respectively, $\eta=-1$), then the feedback $u(t)=-\eta h(t,x(t))$ is deemed to be negative (respectively, positive).
-If negative feedback ensures benign behaviour in the form of convergence of the state to zero, then intuition might suggest that positive feedback causes malign behaviour in the form of non-convergence of the state to zero.
-This intuition finds voice in, for example, [1]: Feedback can upregulate or downregulate a process.
-We refer to upregulation as positive feedback and downregulation as negative feedback.
-Positive feedback … acts as a destabilizing mechanism for the process.
-Negative feedback … acts as a stabilizing mechanism for the process.
-We show, by a simple example, that this positive/negative feedback dichotomy can be invalid.
-In particular, we construct an example (with arbitrary $f$ and $p$) which, under either negative or positive feedback, ensures convergence of the state to zero whilst maintaining boundedness of the control input.
-2 Feedback function
-Let $\varphi\colon\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ be a continuously differentiable bijection with the property that, for some $c_{\varphi}>0$, $\dot{\varphi}(t)\leq c_{\varphi}\big(1+\varphi(t)\big)$ for all $t\geq 0$ (the simplest example being the identity function with $c_{\varphi}=1$).
-Define the set
-$\mathcal{F}:=\{(t,\xi)\in\mathbb{R}_{\geq 0}\times\mathbb{R}\mid~\varphi(t)|\xi|<1\},$ (2.1)
-the continuous function
-$h\colon\mathcal{F}\to\mathbb{R},~~(t,\xi)\mapsto\alpha(\varphi(t)|\xi|)\varphi(t)\xi,$ (2.2)
-$\text{with}~\alpha\colon[0,1)\to[1,\infty),~s\mapsto\frac{1}{1-s},$
-and the feedback-controlled initial-value problem, on the domain $\mathcal{F}$,
-$\dot{x}(t)=f(p(t),x(t))+g(-\eta h(t,x(t)),~~x(0)=x^{0}$ (2.3)
-with $p\in\mathcal{C}(\mathbb{R}_{\geq 0},\mathbb{R})$.
-By a solution we mean a continuous function $x$ on a left-closed interval II (with left endpoint 0) and $\text{graph}(x)\subset\mathcal{F}$, satisfying (2.3).
-A solution is maximal if it has no proper right extension that is also a solution.
-By the standard theory of ordinary differential equations, for each $x^{0}\in\mathbb{R}$, there exists a solution, and every solution has a maximal extension; moreover, the closure of graph of a maximal solution is not a compact subset of $\mathcal{F}$.
-[…]
-Figure 1: Domain $\mathcal{F}$.
-Loosely speaking, the role of the feedback is to maintain the system’s evolution away from the boundary of the domain $\mathcal{F}$ (wherein singularity resides), in which case the evolution continues indefinitely to the right, with transient and asymptotic behaviour determined by the choice of $\varphi$.
-3 Efficacy of both negative and positive feedback
-Let $g$ be given by
-$g\colon\mathbb{R}\to\mathbb{R},~~v\mapsto v\sin\big(\ln(1+|v|)\big).$ (3.1)
-The following technicality will play a central role.
+$\dot{x}(t)=G(p(t),x(t),u(t)),\quad x(0)=x^{0}\in\mathbb{R},$ (1.1)
+where $G\colon\mathbb{R}^{3}\to\mathbb{R}$ and $p\colon\mathbb{R}_{\geq 0}\to\mathbb{R}$ are continuous. In the unperturbed linear case,
+$\dot{x}(t)=ax(t)+bu(t),\quad x(0)=x^{0},\quad a,b\in\mathbb{R},~b\neq 0,$ (1.2)
+the system implicitly assigns (via the parameter $b\neq 0$) a control direction to the input, namely, $\text{\rm{sgn}}(b)$.
+A linear feedback $u(t)=kx(t)$, with $k\neq 0$, is deemed to be positive (respectively, negative) if $\text{\rm{sgn}}(k)=\text{\rm{sgn}}(b)$ (respectively, $\text{\rm{sgn}}(k)=-\text{\rm{sgn}}(b)$).
+In words, the feedback is positive if the polarity of the feedback gain coincides with the control direction (and so $bu(t)=+|bk|x(t)$) and is negative if the polarity of the gain is opposite to the control direction (and so u(t)=|bk|x(t)u(t)=-|bk|x(t)).
+Whilst the concept of positive/negative feedback is unambiguous in the linear case, it is less clear in a nonlinear setting.
+We seek a nonlinear generalization.
+There is a surfeit of candidates for this role.
+We focus on one such candidate based on the observation that, for every compact $K\subset\mathbb{R}$ and $v^{*}\in(0,1)$, the function
+$\chi\colon\mathbb{R}\to\mathbb{R},\quad s\mapsto\min\{v(az+bsv)|~z\in K,~v^{*}\leq|v|\leq 1\}$
+is such that
+$\sup_{s\geq 0}\chi(\eta s)=+\infty\iff\\ \eta~\text{coincides with the control direction}~\text{\rm{sgn}}(b).$
+Extrapolating this observation to the nonlinear context of (1.1), we adopt the following definition of control direction.
+Definition 1.1.
+$\eta\in\{-1,+1\}$ is a control direction for (1.1) if, for every compact $K\subset\mathbb{R}\times\mathbb{R}$, there exists $v^{*}\in(0,1)$ such that the function
+$\chi\colon\mathbb{R}\to\mathbb{R},\\ s\mapsto\min\{vG(\rho,\xi,sv)|~(\rho,\xi)\in K,~v^{*}\leq|v|\leq 1\}$ (1.3)
+has the property $\sup_{s\geq 0}\chi(\eta s)=+\infty$.
+There are three possibilities: system (1.1) may fail to have a control direction (as is the case when $G$ is bounded); (1.1) has unique control direction (as is so when $G$ is linear); (1.1) has non-unique control direction (in Section 3 below, we construct one such example for which each of $\eta=-1$ and $\eta=+1$ qualifies as a control direction).
+For later use, we record the following.
 Proposition 1.2.
-Let $f\colon\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ be continuous, $g$ as in (3.1), and $K,P\subset\mathbb{R}$ compact. Define $V:=[-1,-\tfrac{1}{2}]\cup[\tfrac{1}{2},1]$. The function
-$\chi\colon$ $\mathbb{R}\to\mathbb{R},$
-$s\mapsto\min\{vf(\rho,\xi)+vg(sv)\mid\rho\in P,~\xi\in K,~v\in V\}$
-is continuous and is such that, for each $\eta\in\{-1,+1\}$, $\sup_{s\geq 0}\chi(\eta s)=\infty$.
+Let $G\colon\mathbb{R}^{3}\to\mathbb{R}$ be continuous, $K\subset\mathbb{R}\times\mathbb{R}$ compact and $v^{*}\in(0,1)$. The function $\chi$, given by (1.3), is continuous.
 Proof.
 For notational convenience, introduce the function
-$F\colon\mathbb{R}^{3}\times\mathbb{R}\to\mathbb{R},~~(z,s)=\big((\rho,\xi,v),s\big)\mapsto vf(\rho,\xi)+vg(sv)$
-and write $Z:=P\times K\times V$, a compact set. Observe that
-$\chi\colon s\mapsto\min_{z\in Z}F(z,s)$
+$F\colon\mathbb{R}^{3}\times\mathbb{R}\to\mathbb{R},~~(z,s)=\big((\rho,\xi,v),s\big)\mapsto vG(\rho,\xi,sv).$
+Write $V:=[-1,-v^{*}]\cup[v^{*},1]$ and $Z:=K\times V$, a compact set. Observe that
+$\chi(s)=\min_{z\in Z}F(z,s)~~\forall\,s\in\mathbb{R},$
 and so, for each $s\in\mathbb{R}$, there exists $z_{s}\in Z$ such that $\chi(s)=F(z_{s},s)$.
@@
 In particular, there exist finitely many points $z_{1},\ldots,z_{n}\in Z$ such that
-$Z\times\{s\}\subset\cup_{i=1}^{n}\{(\zeta,\sigma)\mid\|\zeta-z_{i}\|+|\sigma-s|<\delta(z_{i})\}.$
+$Z\times\{s\}\subset\bigcup_{i=1}^{n}\{(\zeta,\sigma)\mid\|\zeta-z_{i}\|+|\sigma-s|<\delta(z_{i})\}.$
 Define $\delta^{*}:=\min\{\delta(z_{1}),\ldots,\delta(z_{n})\}$ and observe that
@@
 $\varepsilon>|F(z_{\sigma},\sigma)-F(z_{\sigma},s)|\geq F(z_{\sigma},s)-F(z_{\sigma},\sigma)\geq\chi(s)-\chi(\sigma).$
-Therefore, $|\chi(\sigma)-\chi(s)|<\varepsilon$ for all $\sigma\in\mathbb{R}$ with $|\sigma-s|<\delta^{*}$, and so $\chi$ is continuous at $s$.
-It remains to show that, for each $\eta\in\{-1,+1\}$, $\sup_{s\geq 0}\chi(\eta s)=+\infty$. Define the increasing and unbounded sequence $(s_{n})$ by
-$s_{n}:=\tfrac{1}{2}e^{(n+1)\pi}-1>0\quad\forall\,n\in\mathbb{N},$
-whence
-$\ln(1+s_{n})=(n+1)\pi-\ln 2\quad\forall\,n\in\mathbb{N}.$
-Also, invoking the inequality $e^{\pi/2}>4$,
-$1+\tfrac{1}{2}s_{n}=\tfrac{1}{2}+\tfrac{1}{4}e^{(n+1)\pi}>e^{n\pi}(\tfrac{1}{4}e^{\pi})>e^{(n+\frac{1}{2})\pi}$
-and so
-$\ln(1+\tfrac{1}{2}s_{n})>n\pi+\tfrac{\pi}{2}\quad\forall\,n\in\mathbb{N}.$
-We may now infer
-$n\pi+\tfrac{\pi}{2}<\ln(1+s_{n}|v|)\leq(n+1)\pi-\ln 2\quad\forall\,v\in V~~\forall\,n\in\mathbb{N}.$
-First consider the case of $n$ even and write $n=2k$. Since the function $\sin$ is decreasing on the interval $\big[\tfrac{\pi}{2},\pi-\ln 2\big]$, for all $v\in V$ and all $k\in\mathbb{N}$ we have
-$1$ $=\sin(2k\pi+\tfrac{\pi}{2})>\sin(\ln(1+s_{2k}|v|))$
-$\geq\sin((2k+1)\pi-\ln 2)=\sin(\pi-\ln 2)=\sin(\ln 2)>0,$
-whence
-$vg(s_{2k}v)=s_{2k}v^{2}\sin(\ln(1+s_{2k}|v|))\geq\tfrac{1}{4}s_{2k}\sin(\ln 2).$
-Therefore, writing
-$c_{1}:=\min\{vf(\rho,\xi)\mid\rho\in P,~\xi\in K,~v\in V\},$
-we have
-$\chi(s_{2k})\geq c_{1}+\min\{vg(s_{2k}v)\mid v\in V\}\geq c_{1}+\tfrac{1}{4}s_{2k}\sin(\ln 2)$
-for all $k\in\mathbb{N}$ and so $\sup_{s\geq 0}\chi(s)=+\infty$.
-Now consider the case of $n$ odd and write n=2k1n=2k-1. Since the function $\sin$ is increasing on the interval $\big[-\tfrac{\pi}{2},-\ln 2\big]$, we have
-$-1$ $=\sin\big((2k-1)\pi+\tfrac{\pi}{2}\big)<\sin(\ln(1+s_{2k-1}|v|))$
-$\leq\sin(2k\pi-\ln 2)=-\sin(\ln 2)\quad\forall\,v\in V,~~\forall\,k\in\mathbb{N},$
-whence
-$vg(-s_{2k-1}v)$ $=-s_{2k-1}v^{2}\sin(\ln(1+s_{2k-1}|v|))$
-$\geq\tfrac{1}{4}s_{2k-1}\sin(\ln 2)\qquad\forall\,v\in V,~~\forall\,k\in\mathbb{N}.$
-Therefore,
-$\chi(-s_{2k-1})>c_{1}+\tfrac{1}{4}s_{2k-1}\sin(\ln 2)\quad\forall\,k\in\mathbb{N}$
-and so $\sup_{s\geq 0}\chi(-s)=+\infty$. ∎
-We arrive at the main focus of the note.
+Therefore, $|\chi(\sigma)-\chi(s)|<\varepsilon$ for all $\sigma\in\mathbb{R}$ with $|\sigma-s|<\delta^{*}$, and so $\chi$ is continuous at $s$. ∎
+Assuming that (1.1) has a control direction $\eta$, the objective is to determine a continuous function $h\colon\mathbb{R}_{\geq 0}\times\mathbb{R}\to\mathbb{R}$ such that, for every $x^{0}\in\mathbb{R}$ and bounded, continuous $p$, application of the feedback $u(t)=-\eta h(t,x(t))$ yields a feedback-controlled initial-value problem which has a solution $x(\cdot)$, every solution has a global extension to a solution on $\mathbb{R}_{\geq 0}$, every global solution is such that $x(t)\to 0$ as $t\to\infty$, and the control input $u\colon t\mapsto-\eta h(t,x(t))$ is bounded.
+If $\eta=+1$ (respectively, $\eta=-1$), then the feedback u(t)=h(t,x(t))u(t)=-h(t,x(t)) (respectively, $u(t)=h(t,x(t))$) is deemed to be negative (respectively, positive).
+If negative feedback ensures benign behaviour in the form of convergence of the state to zero, then intuition might suggest that positive feedback causes malign behaviour in the form of non-convergence of the state to zero.
+There are many circumstances wherein such a benign/malign dichotomy is correct, as discussed in, inter alia, [1].
+In Section 3, we show, by a simple example, that there are also circumstances wherein the dichotomy fails to hold: a feedback and its reversed-polarity counterpart can both be benign.
+2 Feedback system
+Let $\varphi\colon\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ be a continuously differentiable bijection with the property that, for some $c_{\varphi}>0$, $\dot{\varphi}(t)\leq c_{\varphi}\big(1+\varphi(t)\big)$ for all $t\geq 0$ (the simplest example being the identity function with $c_{\varphi}=1$).
+Define the set
+$\mathcal{F}:=\{(t,\xi)\in\mathbb{R}_{\geq 0}\times\mathbb{R}\mid~\varphi(t)|\xi|<1\},$ (2.1)
+the continuous function
+$h\colon\mathcal{F}\to\mathbb{R},~~(t,\xi)\mapsto\alpha(\varphi(t)|\xi|)\varphi(t)\xi,$ (2.2)
+$\text{with}~\alpha\colon[0,1)\to[1,\infty),~s\mapsto\frac{1}{1-s},$
+and the feedback-controlled initial-value problem, on the domain $\mathcal{F}$,
+$\dot{x}(t)=G(p(t),x(t),-\eta h(t,x(t)),~~x(0)=x^{0},$ (2.3)
+with continuous $p:\mathbb{R}_{\geq 0}\to\mathbb{R}$.
+By a solution we mean a continuous function $x$ on a left-closed interval $I$ (with left endpoint 0) and $\text{graph}(x)\subset\mathcal{F}$, satisfying (2.3).
+A solution is maximal if it has no proper right extension that is also a solution.
+By the standard theory of ordinary differential equations, for each $x^{0}\in\mathbb{R}$, there exists a solution, and every solution has a maximal extension; moreover, the closure of the graph of a maximal solution is not a compact subset of $\mathcal{F}$.
+[…]
+Figure 1: Domain $\mathcal{F}$.
+Loosely speaking, the rôle of the feedback is to maintain the system’s evolution away from the boundary of the domain $\mathcal{F}$ (wherein singularity resides), in which case the evolution continues indefinitely to the right, with transient and asymptotic behaviour determined by the choice of $\varphi$.
+The above feedback is based on a wider body of work in the area of “funnel control” (a terminology consistent with Figure 1). The interested reader is referred to [2] for a review of this area.
 Proposition 2.3.
-Let $f\in\mathcal{C}(\mathbb{R},\mathbb{R})$ be arbitrary, $g$ as in (3.1), and $h$ as in (2.2).
-For each $\eta\in\{-1,+1\}$, every $x^{0}\in\mathbb{R}$, and every bounded $p\in\mathcal{C}(\mathbb{R}_{\geq 0},\mathbb{R})$, (2.3) has a solution and every solution has a global extension.
+Assume that $G\colon\mathbb{R}^{3}\to\mathbb{R}$ is continuous and such that (1.1) has a control direction $\eta\in\{-1,+1\}$.
+Let $h$ be as in (2.2).
+For each $x^{0}\in\mathbb{R}$ and every bounded, continuous $p:\mathbb{R}_{\geq 0}\to\mathbb{R}$, (2.3) has a solution and every solution has a global extension.
 Every global solution $x\colon\mathbb{R}_{\geq 0}\to\mathbb{R}$ is such that $x(t)\to 0$ as $t\to\infty$, and the control function $u\colon t\mapsto-\eta h(t,x(t))$ is bounded.
 Proof.
-Let $\eta\in\{-1,+1\}$, $x^{0}\in\mathbb{R}$ and bounded $p\in\mathcal{C}(\mathbb{R}_{\geq 0},\mathbb{R})$ be arbitrary.
+Let $x^{0}\in\mathbb{R}$ and bounded, continuous $p:\mathbb{R}_{\geq 0}\to\mathbb{R}$ be arbitrary.
 Let compact $P\subset\mathbb{R}$ be such that $p(t)\in P$ for all $t\geq 0$.
@@
 Let $x\colon[0,\omega)\to\mathbb{R}$ be a maximal solution.
-It suffices to show existence of  $\varepsilon>0$ such that $\varphi(t)|x(t)|<1-\varepsilon$ for all $t\in[0,\omega)$, in which case it immediately follows that $\omega=\infty$ (as, otherwise, $\text{graph}(x)$ has compact closure in $\mathcal{F}$ which is impossible), $x(t)\to 0$ as $t\to\infty$ (since $\varphi$ is unbounded), and $|u(t)|=|h(t,x(t))|\leq(1-\varepsilon)\alpha(1-\varepsilon)<\infty$ for all $t\geq 0$.
+It suffices to show existence of  $\varepsilon>0$ such that $\varphi(t)|x(t)|<1-\varepsilon$ for all $t\in[0,\omega)$, in which case it immediately follows that $\omega=\infty$ (as, otherwise, $\text{graph}(x)$ has compact closure in $\mathcal{F}$ which is impossible), $x(t)\to 0$ as $t\to\infty$ (since the bijection $\varphi$ is unbounded), and $|u(t)|=|h(t,x(t))|\leq(1-\varepsilon)\alpha(1-\varepsilon)<\infty$ for all $t\geq 0$.
 The task of establishing the existence of such  $\varepsilon>0$ is, in turn, equivalent to proving boundedness of the function $k\colon t\mapsto\alpha(\varphi(t)|x(t)|)$.
@@
 $|x(t)|\leq\frac{1}{\varphi_{\mu}(t)}\quad\forall\,t\in[0,\omega)$
-and so, a fortiori, $x(t)\in K:=[-\mu,\mu]$ for all $t\in[0,\omega)$. As before, define
-$\chi\colon$ $\mathbb{R}_{\geq 0}\to\mathbb{R},$
-$s\mapsto\min\{vf(\rho,\xi)+vg(sv)\mid\rho\in P,~\xi\in K,~v\in V\},$
-$\qquad\qquad\qquad\qquad\qquad\qquad V:=[-1,-\tfrac{1}{2}]\cup[\tfrac{1}{2},1].$
+and so, a fortiori, $x(t)\in M:=[-\mu,\mu]$ for all $t\in[0,\omega)$. Write $K:=P\times M$, $V:=\big[-1,-\tfrac{1}{2}\big]\cup\big[\tfrac{1}{2},1\big]$ and, as before, define
+$\chi\colon\mathbb{R}\to\mathbb{R},~~s\mapsto\min\{vG(\rho,\xi,sv)\mid(\rho,\xi)\in K,~v\in V\}.$
 Let $\nu$ be the function
-$\nu\colon~~\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0},~~s\mapsto\nu(s):=\max\{\chi(\eta s),0\}.$
-By Proposition 3.1, $\chi$ is continuous with $\sup_{s\geq 0}\chi(\eta s)=+\infty$ and so $\nu$ is continuous and surjective.
-Fix $\kappa_{0}>\alpha(1/2)$ arbitrarily and define an increasing sequence $(\kappa_{n})$ in $(\kappa_{0},\infty)$ as follows: choose $\kappa_{1}>\kappa_{0}$ such that $\nu(\kappa_{1})=\kappa_{0}+1$ and set
+$\nu\colon~~\mathbb{R}\to\mathbb{R}_{\geq 0},~~s\mapsto\nu(s):=\max\{\chi(\eta s),0\}.$
+By Proposition 1.2, $\chi$ is continuous with, by the assumption that $\eta$ is a control direction, $\sup_{s\geq 0}\chi(\eta s)=+\infty$ and so $\nu$ is continuous and surjective.
+Fix $\kappa_{0}>\alpha(1/2)$ arbitrarily.
+Define an increasing sequence $(\kappa_{n})$ in $(\kappa_{0},\infty)$ as follows: choose $\kappa_{1}>\kappa_{0}$ such that $\nu(\kappa_{1})=\kappa_{0}+1$ and set
 $\kappa_{n}:=\inf\{\kappa>\kappa_{n-1}\mid\nu(\kappa)=\kappa_{0}+n\}\quad\forall\,n\geq 2.$
@@
 Figure 3: Schematic construction of the sequence $(\sigma_{n})$.
-Now define the sequence $(\sigma_{n})$ in $(0,\omega)$ by
+Next, define the sequence $(\sigma_{n})$ in $(0,\omega)$ by
 $\sigma_{n}:=\sup\{t\in[\tau_{n},\tau_{n+1}]\,\mid\nu(k(t))=\kappa_{0}+n\},\quad n\in\mathbb{N}.$
@@
 $\alpha(\varphi(t)|x(t)|)=k(t)<\alpha(1/2)<\kappa_{0}<\kappa_{n}<\kappa_{n+1}=k(\tau_{n+1})$
-Therefore, there exists $t^{*}\in(\sigma_{n},\tau_{n+1})$ such that $k(t^{*})=\kappa_{n}$ whence the contradiction
+Therefore, there exists $t^{*}\in(\sigma_{n},\tau_{n+1})$ such that $k(t^{*})=\kappa_{n}$ and we arrive at the contradiction
 $\kappa_{0}+n<\nu(k(t^{*}))=\nu(\kappa_{n})=\kappa_{0}+n.$
@@
 By symmetry of $V$, we know that $\varphi(t)x(t)\in V$ if, and only if, $\varphi(t)x(t)\in V$, and so we may infer that
-$-\varphi(t)x(t)f(p(t),x(t))-\varphi(t)x(t)g(-\eta k(t)\varphi(t)x(t))$
-$\ \geq\min\{vf(\rho,\xi)+vg(\eta k(t)v)\mid\rho\in P,~\xi\in K,~v\in V\}$
-$\ =\chi(\eta k(t))\quad\forall\,t\in[\sigma_{n},\tau_{n+1}]\quad\forall\,n\in\mathbb{N}$
+$-\varphi(t)x(t)$ $G\big(p(t),x(t),-\eta k(t)\varphi(t)x(t)\big)$
+$\ \geq\min\{vG(\rho,\xi,\eta k(t)v)\mid(\rho,\xi)\in K,~v\in V\}$
+$\ =\chi(\eta k(t))\quad\forall\,t\in[\sigma_{n},\tau_{n+1}]\quad\forall\,n\in\mathbb{N}.$
 Also (noting that, by bijectivity, $\varphi$ is strictly increasing), we have
-$\varphi(t)\dot{\varphi}(t)x(t)^{2}$ $\leq\frac{\dot{\varphi}(t)}{\varphi(t)}\leq c_{\varphi}\big((1/\varphi(t))+1\big)$
+$\varphi(t)\dot{\varphi}(t)x^{2}(t)$ $\leq\frac{\dot{\varphi}(t)}{\varphi(t)}\leq c_{\varphi}\big((1/\varphi(t))+1\big)$
 $\leq c_{\varphi}\big((1/\varphi(\tau_{1}))+1\big)=:\gamma\qquad\forall\,t\in[\tau_{1},\omega).$
 We may now deduce that
-$\tfrac{1}{2}\tfrac{\mathrm{d}}{\mathrm{d}t}(\varphi(t)x(t))^{2}$ $=\varphi(t)\dot{\varphi}(t)x(t)^{2}$
-$\quad+\varphi(t)^{2}x(t)\big(f(p(t),x(t))+g(-\eta h(t,x(t))\big)$
-$\leq\gamma-\varphi(t)\big(-\varphi(t)x(t)f(p(t),x(t))$
-$\quad-\varphi(t)x(t)g(-\eta k(t)\varphi(t)x(t))\big)$
+$\tfrac{1}{2}\tfrac{\mathrm{d}}{\mathrm{d}t}$ $(\varphi(t)x(t))^{2}=\varphi(t)\dot{\varphi}(t)x^{2}(t)$
+$\qquad\qquad\qquad\quad+\varphi^{2}(t)x(t)G\big(p(t),x(t),-\eta h(t,x(t)\big)$
+$\leq\gamma-\varphi(t)\big(-\varphi(t)x(t)G\big(p(t),x(t),-\eta k(t)\varphi(t)x(t)\big)\big)$
 $\leq\gamma-\varphi(t)\chi(\eta k(t))$
@@
 $\big(\varphi(\tau_{n+1})x(\tau_{n+1})\big)^{2}<\big(\varphi(\sigma_{n})x(\sigma_{n})\big)^{2}$
-and we arrive at a contradiction
+and we arrive at the contradiction
 $k(\sigma_{n})$ $<k(\tau_{n+1})=\alpha(\varphi(\tau_{n+1})|x(\tau_{n+1})|)$
@@
 Therefore, the supposition of unboundedness of $k(\cdot)$ is false. ∎
+The following corollary is immediate, and constitutes the essence of the note.
+Corollary 2.4.
+Let $h$ be as in (2.2).
+If system (1.1) has non-unique control direction, then the control objective is achieved by both the feedback u(t)=h(t,x(t))u(t)=-h(t,x(t)) and its polarity-reversed counterpart $u(t)=h(t,x(t))$.
+3 Example: efficacy of both negative and positive feedback
+Finally, we proceed to construct an example of a system of form (1.1) with non-unique control direction.
+This example, in conjunction with Corollary 2.4, establishes the efficacy of a feedback and its polarity-reversed counterpart.
+Let $G$ be given by
+$G\colon(\rho,\xi,v)\mapsto f(\rho,\xi)+bg(v),\quad b\neq 0,$ (3.1)
+with $f\colon\mathbb{R}\to\mathbb{R}$ continuous and
+$g\colon\mathbb{R}\to\mathbb{R},~~v\mapsto v\sin\big(\ln(1+|v|)\big).$ (3.2)
+Proposition 3.5.
+System (1.1), with $G$ given by (3.1), has non-unique control direction.
+Proof.
+Let $K\subset\mathbb{R}\times\mathbb{R}$ be compact and $V:=\big[-1,-\tfrac{1}{2}\big]\cup\big[\tfrac{1}{2},1\big]$. We will show that, for each $\eta\in\{-1,+1\}$,
+$\sup_{s\geq 0}\chi(\eta s)=+\infty,$
+where $\chi\colon s\mapsto\min\{vf(\rho,\xi)+bvg(sv)\mid(\rho,\xi)\in K,~v\in V\}.$
+Define the increasing and unbounded sequence $(s_{n})$ by
+$s_{n}:=\tfrac{1}{2}e^{(n+1)\pi}-1>0\quad\forall\,n\in\mathbb{N},$
+and so
+$\ln(1+s_{n})=(n+1)\pi-\ln 2\quad\forall\,n\in\mathbb{N}.$
+Also, invoking the inequality $e^{\pi/2}>4$,
+$1+\tfrac{1}{2}s_{n}=\tfrac{1}{2}+\tfrac{1}{4}e^{(n+1)\pi}>e^{n\pi}(\tfrac{1}{4}e^{\pi})>e^{(n+\frac{1}{2})\pi}$
+and so
+$\ln(1+\tfrac{1}{2}s_{n})>n\pi+\tfrac{\pi}{2}\quad\forall\,n\in\mathbb{N}.$
+We may now infer
+$n\pi+\tfrac{\pi}{2}<\ln(1+s_{n}|v|)\leq(n+1)\pi-\ln 2\quad\forall\,v\in V~~\forall\,n\in\mathbb{N}.$
+First consider the case of $n$ even and write $n=2k$. Since the function $\sin$ is decreasing on the interval $\big[\tfrac{\pi}{2},\pi-\ln 2\big]$, for all $v\in V$ and all $k\in\mathbb{N}$ we have
+$1$ $=\sin(2k\pi+\tfrac{\pi}{2})>\sin(\ln(1+s_{2k}|v|))$
+$\geq\sin((2k+1)\pi-\ln 2)=\sin(\pi-\ln 2)=\sin(\ln 2)>0.$
+Writing $\sigma_{k}:=\text{\rm{sgn}}(b)s_{k}$ for all $k\in\mathbb{N}$, we find
+$bvg(\sigma_{2k}v)$ $=|b|\text{\rm{sgn}}(b)\,vg(s_{2k}\text{\rm{sgn}}(b)v)$
+$=s_{2k}|b|v^{2}\sin(\ln(1+s_{2k}|v|))\geq\tfrac{1}{4}s_{2k}|b|\sin(\ln 2).$
+Therefore, writing
+$c_{1}:=\min\{vf(\rho,\xi)\mid(\rho,\xi)\in K,~v\in V\},$
+we have
+$\chi(\sigma_{2k})$ $\geq c_{1}+\min\{bvg(\sigma_{2k}v)\mid v\in V\}$
+$\geq c_{1}+\tfrac{1}{4}s_{2k}|b|\sin(\ln 2)\quad\forall\,k\in\mathbb{N}.$
+For $\eta=\text{\rm{sgn}}(b)$, it follows that $\chi(\eta s_{2k})\to+\infty$ as $k\to\infty$ and so $\sup_{s\geq 0}\chi(\eta s)=+\infty$.
+Now consider the case of $n$ odd and write n=2k1n=2k-1. The function $\sin$ is increasing on the interval $\big[-\tfrac{\pi}{2},-\ln 2\big]$ and so
+$-1$ $=\sin\big((2k-1)\pi+\tfrac{\pi}{2}\big)<\sin(\ln(1+s_{2k-1}|v|))$
+$\leq\sin(2k\pi-\ln 2)=-\sin(\ln 2)\quad\forall\,v\in V~~\forall\,k\in\mathbb{N}.$
+Therefore,
+$bvg(-\sigma_{2k-1}v)$ $=-|b|s_{2k-1}v^{2}\sin(\ln(1+s_{2k-1}|v|))$
+$\geq\tfrac{1}{4}|b|s_{2k-1}\sin(\ln 2)\quad\forall\,v\in V~~\forall\,k\in\mathbb{N}$
+and we have
+$\chi(-\sigma_{2k-1})\geq c_{1}+\tfrac{1}{4}s_{2k-1}\sin(\ln 2)\quad\forall\,k\in\mathbb{N}.$
+For $\eta=-\text{\rm{sgn}}(b)$, it follows that $\chi(\eta s_{2k-1})\to\infty$ as $k\to\infty$, and so $\sup_{s\geq 0}\chi(\eta s)=+\infty$.
+In summary, it has been shown that, for each $\eta\in\{-1,+1\}$, $\sup_{s\geq 0}\chi(\eta s)=+\infty$. ∎
+4 Discussion
+In succinctly resolving a conjecture by Morse in the 1980s, Nussbaum [3] sparked activity in an area now often referred to as universal adaptive control.
+In the simplest of contexts, namely, that of scalar linear systems of form (1.2), a basic question is: “in the absence of knowledge of the control direction $\text{\rm{sgn}}(b)$, does there exist a (reasonably regular) feedback strategy which ensures that the zero state is globally attractive?”.
+As is now well known, this question can be answered in the affirmative: the adaptive strategy
+$u(t)=N(k(t))x(t),\quad\dot{k}(t)=x(t)^{2},\quad k(0)=k^{0}\geq 0$
+achieves the control objective if the function $N$ has the properties:
+$\sup_{\kappa>0}\,\frac{1}{\kappa}\int_{0}^{\kappa}N(k)\textrm{d}k=+\infty\,;\quad\inf_{\kappa>0}\,\frac{1}{\kappa}\int_{0}^{\kappa}N(k)\textrm{d}k=-\infty.$
+For example, the function $N\colon\kappa\mapsto\kappa\sin\sqrt{\kappa}$ suffices.
+Loosely speaking, $N$ ensures that the control gain $N\circ k$ can “probe” in each control direction and has the potential to take arbitrarily large values in magnitude.
+Conceptually, the function $g$ in (3.1) plays a similar role.
+However, the comparison diverges.
+In the adaptive context, the “probing” mechanism is a design feature of the controller, whereas, in the context of the present note, the potential for “probing” is implicit in the system itself: moreover, the adaptive control structure is dynamic, whereas the control structure (2.2) is static (albeit time-dependent).
 5 Conclusion
-The claim in this note is purely existential in nature in the sense that a commonly-perceived dichotomy of negative and positive feedback is shown not to be valid in general.
-This has been done by proving the existence of a nonlinear counter-example: the practical utility and quantitative behaviour of the feedback employed therein has not been addressed.
-Whilst the counter-example is open to a criticism of artificiality, it does, however, serve to illustrate that the dichotomy “negative/positive feedback” equates to “good/bad behaviour” may be inaccurate.
-The counter-example is based on a wider body of work in the area of “funnel control” (a terminology consistent with Figure 1).
-The interested reader is referred to [2] for a review of this area.
+A dichotomy of negative and positive feedback has been shown to be invalid in general.
+This has been done by establishing the existence of a nonlinear example which serves to illustrate that “negative/positive feedback” need not equate to “good/bad behaviour”.
+In this sense, the claim in this note is purely existential in nature.
+In essence, it simply provides an example in which a feedback and its reversed-polarity counterpart both cause “good” behaviour.
+The practical utility and quantitative behaviour of the feedback employed therein have not been addressed.
+The particular example in Section 3 may be open to a criticism of artificiality: we stress that its role is only to show that a system of the general form (1.1) can have non-unique control direction, thereby ensuring the existence of a stabilizing feedback with the property that its reversed-polarity counterpart is also stabilizing.

### v3

@@
 There are three possibilities: system (1.1) may fail to have a control direction (as is the case when $G$ is bounded); (1.1) has unique control direction (as is so when $G$ is linear); (1.1) has non-unique control direction (in Section 3 below, we construct one such example for which each of $\eta=-1$ and $\eta=+1$ qualifies as a control direction).
-For later use, we record the following.
+In essence, Definition 1.1 attempts to capture the following intuitive idea (loosely stated).
+The existence of a control direction is a system requirement that large input values $v$ should dominate in the sense that, with the pair $(\rho,\xi)$ of its first two arguments compactly constrained, the function $G$ is such that $|G(\rho,\xi,v)|$ can attain arbitrarily large values by taking $|v|$ sufficiently large.
+For example, if $G\colon(\rho,\xi,v)\mapsto f(\rho,\xi)g(v)$, then surjectivity of (continuous) $g$ is a necessary condition for the existence of a control direction.
+For later use, we record the following fact (a particular consequence of a general result in variational analysis ).
 Proposition 1.2.
 Let $G\colon\mathbb{R}^{3}\to\mathbb{R}$ be continuous, $K\subset\mathbb{R}\times\mathbb{R}$ compact and $v^{*}\in(0,1)$. The function $\chi$, given by (1.3), is continuous.
-Proof.
-For notational convenience, introduce the function
-$F\colon\mathbb{R}^{3}\times\mathbb{R}\to\mathbb{R},~~(z,s)=\big((\rho,\xi,v),s\big)\mapsto vG(\rho,\xi,sv).$
-Write $V:=[-1,-v^{*}]\cup[v^{*},1]$ and $Z:=K\times V$, a compact set. Observe that
-$\chi(s)=\min_{z\in Z}F(z,s)~~\forall\,s\in\mathbb{R},$
-and so, for each $s\in\mathbb{R}$, there exists $z_{s}\in Z$ such that $\chi(s)=F(z_{s},s)$.
-Fix $s\in\mathbb{R}$ arbitrarily.
-We will show that $\chi$ is continuous at $s$.
-Let $\varepsilon>0$.
-By continuity of $F$, for each $z\in Z$, there exists $\delta(z)>0$ such that, for all $\zeta,\sigma\in\mathbb{R}$ we have
-$\|\zeta-z\|+|\sigma-s|<\delta(z)\implies|F(\zeta,\sigma)-F(z,s)|<\varepsilon$
-and so $\cup_{z\in Z}\{(\zeta,\sigma)\mid\|\zeta-z\|+|\sigma-s|<\delta(z)\}$ is an open cover of $Z\times\{s\}$.
-By compactness of the latter, there exists a finite sub-cover.
-In particular, there exist finitely many points $z_{1},\ldots,z_{n}\in Z$ such that
-$Z\times\{s\}\subset\bigcup_{i=1}^{n}\{(\zeta,\sigma)\mid\|\zeta-z_{i}\|+|\sigma-s|<\delta(z_{i})\}.$
-Define $\delta^{*}:=\min\{\delta(z_{1}),\ldots,\delta(z_{n})\}$ and observe that
-$\forall\,\zeta,\sigma\in\mathbb{R}\colon~~z\in Z,~|\sigma-s|<\delta^{*}\implies|F(z,\sigma)-F(z,s)|<\varepsilon.$
-It follows that, for all $\sigma\in\mathbb{R}$ with $|\sigma-s|<\delta^{*}$ we have
-$\varepsilon>|F(z_{s},\sigma)-F(z_{s},s)|\geq F(z_{s},\sigma)-F(z_{s},s)\geq\chi(\sigma)-\chi(s)$
-and, similarly,
-$\varepsilon>|F(z_{\sigma},\sigma)-F(z_{\sigma},s)|\geq F(z_{\sigma},s)-F(z_{\sigma},\sigma)\geq\chi(s)-\chi(\sigma).$
-Therefore, $|\chi(\sigma)-\chi(s)|<\varepsilon$ for all $\sigma\in\mathbb{R}$ with $|\sigma-s|<\delta^{*}$, and so $\chi$ is continuous at $s$. ∎
 Assuming that (1.1) has a control direction $\eta$, the objective is to determine a continuous function $h\colon\mathbb{R}_{\geq 0}\times\mathbb{R}\to\mathbb{R}$ such that, for every $x^{0}\in\mathbb{R}$ and bounded, continuous $p$, application of the feedback $u(t)=-\eta h(t,x(t))$ yields a feedback-controlled initial-value problem which has a solution $x(\cdot)$, every solution has a global extension to a solution on $\mathbb{R}_{\geq 0}$, every global solution is such that $x(t)\to 0$ as $t\to\infty$, and the control input $u\colon t\mapsto-\eta h(t,x(t))$ is bounded.